OpenStudy (anonymous):

How do i find the inverse? I need to find the inverse of g(x) = 2/3x - 4

OpenStudy (anonymous):

switch x & y and solve for y

OpenStudy (mertsj):

Write y = 2/3x-4 Switch the y and x and solve the resulting equation for y. That will be the inverse.

OpenStudy (anonymous):

like x=2/3y-4 <-- here solve for y

OpenStudy (anonymous):

so whats the answer, Is it x+4/ 2/3 = y

OpenStudy (anonymous):

I dont get it

OpenStudy (anonymous):

so its, x=2/3y-4 x+4=2/3y 3/2(x+4)=y

OpenStudy (anonymous):

that the answer

OpenStudy (anonymous):

yup..